We have a domain where there is only one feature, X, which is binary. The domain has also a category label Y, which is binary as well.

Given a Naive Bayes model with following parameters:

P(X=T|Y=T) = 0.2

P(X=F|Y=F) = 0.6

P(Y = T) = 0.7

Given an instance X = T, what is the probability that its label is True calculated by the model?
Answer #1

here P(X=T)=P(Y=T)*P(X=T|Y=T)+P(Y=F)*P(X=T|Y=F)=0.7*0.2+(1-0.7)*(1-0.6)=0.14+0.12=0.26

hence P(Y=T|X=T)=P(Y=T)*P(X=T|Y=T)/P(X=T)=0.7*0.2/0.26=0.538462